# Latent Diffusion Prior Enhanced Deep Unfolding for Snapshot Spectral Compressive Imaging

### v1

@@ 3.3 Latent Diffusion Prior Assisted Unfolding Denoising
The denoising process in deep unfolding leads to a natural performance bottleneck due to the intrinsic problem of heavily degraded input. Thus, we introduce external degradation-free prior knowledge to adaptively compensate for the denoising process. We will then introduce this process in a two-phase manner.
Phase I: Learning Prior Knowledge from clean HSIs. In this phase, we use an image encoder to compress both compressive measurement $\bm{y}$ and clean HSIs (Ground-Truth hyperspectral images) $\bm{x}$ into latent space. However, instead of simply using measurement $\bm{y}$, we transfer $\bm{y}$ by Euclidean projection to 3D HSIs space and normalize it by sensing matrix $\bm{y}_{norm}\in\mathbb{R}^{W\times H\times N_{\lambda}}=\bm{A}^{\top}\left(\bm{A}\bm{A}^{\top}\right)^{-1}\bm{y}$. This will improve the balance between two different inputs and easier for the encoder to learn their relation. The input of the encoder in the first phase is ${\bm{I}}\in\mathbb{R}^{W\times H\times 2B}=\text{concatenate}(\bm{y}_{norm},\bm{x})$. Thus the latent encoder process can be written as $\bm{z}_{GT}\in\mathbb{R}^{N\times C}=\text{LE}({\bm{I}})$, where $N\ll W\times H$, $C$ is the latent feature channel number. The LE can be seen in Fig. 3 (c), it consists of several residual convolution and linear operations. Then this learned representation $\bm{z}_{GT}$ will be used as prior in the denoiser to compensate for the denoising errors. The Deep Unfolding Network (DUN) will reconstruct HSI signals using measurement and mask with the assistance of $\bm{z}_{GT}$, i.e. $\hat{\bm{x}}=\text{DUN}(\bm{y},\bm{A},\bm{z}_{GT}).$ In this phase, we only use the reconstruction loss: $\mathcal{L}_{\text{rec}}=\left\|\bm{x}-\hat{\bm{x}}\right\|_{1}$.
Phase II: Generating Prior by Latent Diffusion Model. After learning the prior representation from clean HSIs, we aim to learn an LDM to generate this prior condition on measurement $\bm{y}$ in the second phase. Specifically, the encoder in the first phase LE is fixed to encode clean HSIs and measurements to $\bm{z}$ as the generative object of latent space, i.e. the starting point of the forward Markov process in the diffusion model. Then as usual forward process, Gaussian noise will be gradually added on $\bm{z}$ across $T$ time steps according to the parameter $\beta_{t}$, stated as:
@@
where $\bm{z}_{t}$ represents the noisy features at the $t$-th step, and $\bm{z}_{0}=\bm{z}_{GT}$ is the generative target. $\beta_{1:T}\in(0,1)$ are hyperparameters that control the variance of the Gaussian distribution $\mathcal{N}$. Through iterative derivation with reparameterization [25], Eq. equation 8 can be written as:
@@
where $\bm{\epsilon}$ represents the noise added on $\bm{z}_{t}$, which is the only uncertain variable. Thus, we adopt a denoising network, denoted as $\bm{\epsilon}_{\theta}$, to estimate the noise $\bm{\epsilon}$ at each step, following the previous works [18, 46, 44]. Considering our diffusion model operates in the latent space, we utilize another latent encoder to extract latent features, denoted as $\text{LE}^{\prime}$, with the same structure as the latent encoder of phase I. Specifically, $\text{LE}^{\prime}$ compresses the normalized measurement $\bm{y}_{norm}$ into latent space to get the condition latent $\mathbf{c}\in\mathbb{R}^{N\times C^{\prime}}$. In the end, we use the denoising network to predict the noise $\epsilon_{t}$ according to $\bm{z}_{t}$ of the previous step in reverse process and the condition features $\mathbf{c}$, stated as $\bm{\epsilon}=\bm{\epsilon}_{\theta}\left(\bm{z}_{t},\mathbf{c},t\right)$. With the substitution of $\bm{\epsilon}_{\theta}$ in Eq. (10) and set the variance as $1-\alpha_{t}$, the reverse inference can be stated as:
$\bm{z}_{t-1}=\textstyle\frac{1}{\sqrt{\alpha_{t}}}\left(\bm{z}_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\bm{\epsilon}_{\theta}\left(\bm{z}_{t},\mathbf{c},t\right)\right)+\sqrt{1-\alpha_{t}}\bm{\epsilon}_{t},$ (11)
where $\bm{\epsilon}_{t}\sim\mathcal{N}(0,\mathbf{I})$. Finally, we can generate the target prior feature $\hat{\bm{z}}\in\mathbb{R}^{N\times C^{\prime}}$ after $T$ iterative sampling $\bm{z}_{t}$ by Eq. (11). As shown in Fig. 3(d), the predicted prior feature is then used to guide the Transformer in denoiser. Notably, since the distribution of the latent space with the size of $\mathbb{R}^{N\times C}$ (e.g, $16\times 256$) is much simpler than that of images with size $\mathbb{R}^{H\times W\times B}$ (e.g, $256\times 256\times 28$), the prior feature can be generated with a small number of iterations $T$, corresponding to paper [44].
@@
$\nabla_{\bm{\theta}}\left\|\bm{\epsilon}-\bm{\epsilon}_{\theta}\left(\sqrt{\bar{\alpha}_{t}}\bm{z}+\sqrt{1-\bar{\alpha}_{t}}\bm{\epsilon},\mathbf{c},t\right)\right\|_{2}^{2},$ (12)
where $\bm{z}$ and $\mathbf{c}$ are ground-truth prior features and the latent condition representations defined above; $t\in[1,T]$ is the randomly sampled time step; $\bm{\epsilon}\sim\mathcal{N}(0,\mathbf{I})$ denotes the sampled Gaussian noise. We jointly update all the parameters in the network with the objective loss function of the second phase, including: the deep unfolding network, the feature encoder $\text{LE}^{\prime}$, and the diffusion denoising network $\bm{\epsilon}_{\theta}$. The objective loss function of the second phase can be stated as:
@@ 3.4 Aggregate Features by Trident Transformer
Figure 4: The illustration of (a) the overall structure of Trident Transformer, and (b)-(d) the detailed sub-modules in the Trident Transformer.
@@
Previous HSI reconstruction methods usually only exploit the relation between spatial and spectral, both externally and internally. However, the spatial-spectral relations are challenging to explore only with compressed measurements. Therefore, we design a Transformer, named Trident Transformer (TT), to effectively aggregate our learned high-quality degradation-free prior knowledge for compensation.
@@
Firstly, inspired by the multi-scale operations in previous papers [9, 58] with hierarchical structures, we downsample the prior to obtain the multi-scale prior representations. Specifically, three downsampling layers are employed, and the outputs contain prior features of three scales, stated as:
$\bm{z}^{i}=\left\{\begin{array}[]{ccl}\bm{z}_{GT},&{if}&i=1,\\ \text{downsample}(\bm{z}^{i-1}),&{if}&i>1\end{array}\right.~{},$ (14)
@@
where $\bm{z}^{i}\in\mathbb{R}^{\frac{N}{2^{i}-1}\times 2^{i-1}C},~{}i=1,2,3$. For $i=1$, $\bm{z}^{i}=\bm{z}_{GT}$, which is computed in the first phase training; For $i=2,3$, $\bm{z}^{i}=\hat{\bm{z}}$, which is utilized for training and inference in the second phase.
@@
As shown in Fig. 4, our Trident Transformer includes three branches: spatial flow, cross-spectral flow, and cross-prior flow. Each branch shares the information flow with others and is then fused by the aggregation layer and a feed-forward network (FFN). Before the embedding layer, the input feature at $i$-th scale ${{\bf U}}_{i}\in\mathbb{R}^{H_{i}\times W_{i}\times C_{i}}$ is split into ${{\bf U}}^{C}_{i}\in\mathbb{R}^{H_{i}\times W_{i}\times\frac{C_{i}}{2}}$ and ${{\bf U}}^{S}_{i}\in\mathbb{R}^{H_{i}\times W_{i}\times\frac{C_{i}}{2}}$ along the channel dimension, denoting cross flow input and spatial flow input respectively. The spatial flow consists a series of mobile blocks [21] without batch norms, which aims to extract representative visual information with fewer parameters and FLOPs.
3.4.1 Cross Spectral Flow
@@
In cross spectral flow (CSF) module, as shown in Fig. 4 (c), we design a shrunken multi-head self-attention with cross-scale embedding. This flow primarily focuses on the spectral dimension and aims to save computational burden acoording to the spatial size. Specifically, we compress the spatial size of the query embedding (${{\bf Q}}$) and key embedding (${{\bf K}}$) to $1/4$ and expand its channel twice. After establishing the spectral correlation, we use a dilation convolution on the value embedding (${{\bf V}}$) to obtain larger perceptual field information with expanded channel dimension and unchanged spatial dimension. The CSF can be formulated as:
${MSA_{i}^{CS}}({{\bf U}}_{i})={{\bf W}}_{c1}^{CS}{{\bf V}}_{i}^{CS}\odot\text{Softmax}({{\bf K}}_{i}^{CS}\odot{{\bf Q}}_{i}^{CS}/\alpha),$ (15)
${{\bf Q}}_{i}^{CS}={{\bf W}}^{QCS}{{\bf U}}_{i},~{}~{}{{\bf K}}_{i}^{CS}={{\bf W}}^{KCS}{{\bf U}}_{i},~{}~{}{{\bf V}}_{i}^{CS}={{\bf W}}^{VCS}{{\bf U}}_{i},$ (16)
where ${{\bf W}}^{*}$ represent the weights of bias-free convolution.
3.4.2 Cross Prior Flow
@@
Cross prior flow (CPF) in Fig. 4 (d) is a variable shared multi-head cross-attention. The query in this flow is borrowed from the value of CSF which is extracted from a large perceptive field with more spatial information. In this way, the prior could facilitate to compensate for spatial deficiency. Compared to the spectral recovery, the spatial recovery is typically more challenging. Our manipulation can be formulated as:
${MSA_{i}^{CP}}({{\bf U}}_{i})=W_{c1}{{\bf V}}\odot\text{Softmax}({{\bf K}}\odot{{\bf Q}}/\alpha),$ (17)
@@
3.4.3 Flow Interaction and Aggregation
In order to compensate for the deficiency of spatial information in CSF and CPF, and the spectral information in the spatial flow, we fuse the compensation information together to reconstruct hyperspectra images. As shown in Fig. 4, the colorful arrows represents the information interactions between each flow. Specifically, information of each module modulate with other flows, where the $1\times 1$ convolutions serve as compensation bridges. The aggregation part consists of concatenation, convolution layers, and an activation function for a weighted combination of each flow output. In our Trident Transformer, the prior knowledge learned from the clean images will provide compensation for reconstruction in both spatial and spectral details, avoiding the influence of degraded measurements.
Table 1: The average results of PSNR in dB (top entry in each cell), SSIM (bottom entry in each cell) on the 10 synthetic spectral scenes.
@@ 3.4.3 Flow Interaction and Aggregation
Ours-9stg 38.07 41.16 43.70 48.01 37.76 37.65 38.58 36.31 42.66 35.18 39.91
0.969 0.982 0.983 0.993 0.980 0.980 0.973 0.979 0.984 0.967 0.979
Ours-10stg 38.18 41.29 43.65 47.87 37.94 37.72 38.94 36.40 42.94 35.16 40.01
0.970 0.983 0.983 0.994 0.981 0.980 0.974 0.979 0.985 0.968 0.980
Method PSNR (dB) FLOPs (G) Time (ms/10scenes)
Our Full Model 38.31 37.71 421.96
w/o prior 37.63 36.66 403.90
w/o prior and TT 38.11 31.60 305.55
Table 2: Ablation study of our method. ‘Time’ denotes the total inference time of each method dealing with 10 synthetic test scenes.
Figure 6: The real data comparisons. 2 out of 28 wavelengths are selected for visual comparison.

### v2

@@ 4.2 Latent Diffusion Prior Assisted Unfolding Denoising
The denoising process in DUN leads to a natural performance bottleneck due to the intrinsic problem of heavily degraded input. Thus, we introduce external degradation-free prior knowledge to compensate for the denoising process. We will then introduce this process in a two-phase manner.
Phase I: Learning Prior Knowledge from clean HSIs. In this phase, we use an image encoder to compress both compressive measurement $\bm{y}$ and clean HSIs (Ground-Truth hyperspectral images) $\bm{x}$ into latent space. However, instead of simply using measurement $\bm{y}$, we transfer $\bm{y}$ by Euclidean projection to 3D HSIs space and normalize it by sensing matrix $\bm{y}_{norm}\in\mathbb{R}^{W\times H\times N_{\lambda}}=\bm{A}^{\top}\left(\bm{A}\bm{A}^{\top}\right)^{-1}\bm{y}$. This will improve the balance between two different inputs and easier for the encoder to learn their relation. The input of the encoder in the first phase is ${\bm{I}}_{\text{E}}\in\mathbb{R}^{W\times H\times 2N_{\lambda}}=\text{concatenate}(\bm{y}_{norm},\bm{x})$. Thus the latent encoder process can be written as $\bm{z}_{GT}\in\mathbb{R}^{N\times C}=\text{LE}({\bm{I}}_{\text{E}})$, where $N\ll W\times H$, $C$ is the latent feature channel number. The LE can be seen in Fig. 3 (c): to alleviate the computational burden within LDM, we employ mobile blocks (MBlocks) [24], devoid of batch normalization instead of normal convolution. It aims at efficiently extracting representative visual features while maintaining computational efficiency. Moreover, considering the limitations in convolution, we add an MLP-Mixer [58] in LE to provide fast information exchange between patches by token-mixing MLP. Then the $\bm{z}_{GT}$ will be used as prior in the denoiser to compensate for the denoising errors. The DUN will reconstruct HSI signals using measurement and mask with the assistance of $\bm{z}_{GT}$, i.e. $\hat{\bm{x}}=\text{DUN}(\bm{y},\bm{A},\bm{z}_{GT}).$ Note that unlike the original LDM having its entire ‘Auto Encoder’, there is no corresponding ‘decoder’ specified here, because $\bm{z}$ is sent to DUN for ‘decoding’. In this phase, we only use the reconstruction loss: $\mathcal{L}_{\text{rec}}=\left\|\bm{x}-\hat{\bm{x}}\right\|_{1}$.
Phase II: Generating Prior by Latent Diffusion Model. After learning the prior representation from clean HSIs, we aim to learn an LDM to generate this prior condition on measurement $\bm{y}$ in the second phase. Specifically, the encoder in the first phase LE is fixed to encode clean HSIs and measurements to $\bm{z}_{GT}$ as the generative object of latent space, i.e. the starting point of the forward Markov process in the diffusion model. Then as usual forward process, Gaussian noise will be gradually added on $\bm{z}_{GT}$ across $T$ time steps according to the parameter $\beta_{t}$:
@@
where $\bm{z}_{t}$ represents the noisy features at the $t$-th step, and $\bm{z}_{0}=\bm{z}_{GT}$ is the generative target. $\beta_{1:T}\in(0,1)$ are hyperparameters that control the variance of the Gaussian distribution $\mathcal{N}$. Through iterative derivation with reparameterization [28], Eq. (7) can be written as:
@@
where $\bm{\epsilon}$ represents the noise added on $\bm{z}_{t}$. Thus, a denoising network $\bm{\epsilon}_{\theta}$ is used to predict the noise $\bm{\epsilon}$ at each step, following the previous works [22, 53]. In order to encode condition $\bm{y}$ to latent space, another encoder $\text{LE}^{\prime}$ is applied to extract features, with the same structure as the LE of Phase I. Specifically, $\text{LE}^{\prime}$ compresses the normalized measurement $\bm{y}_{norm}$ into latent space to get the latent condition features ${\bm{c}}\in\mathbb{R}^{N\times C}$. In the end, we use the denoising network to predict the noise $\epsilon_{t}$ according to $\bm{z}_{t}$ of the previous step in reverse process and the condition ${\bm{c}}$, stated as $\bm{\epsilon}=\bm{\epsilon}_{\theta}\left(\bm{z}_{t},{\bm{c}},t\right)$. With the substitution of $\bm{\epsilon}_{\theta}$ in Eq. (9) and set the variance as $1-\alpha_{t}$, the reverse inference can be stated as:
$\bm{z}_{t-1}=\textstyle\frac{1}{\sqrt{\alpha_{t}}}\left(\bm{z}_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\bm{\epsilon}_{\theta}\left(\bm{z}_{t},{\bm{c}},t\right)\right)+\sqrt{1-\alpha_{t}}\bm{\epsilon}_{t},$ (10)
where $\bm{\epsilon}_{t}\sim\mathcal{N}(0,\mathbf{I})$. Finally, we can generate the target prior feature $\hat{\bm{z}}\in\mathbb{R}^{N\times C}$ after $T$ iterative sampling $\bm{z}_{t}$ by Eq. (10). As shown in Fig. 3(e), the predicted prior feature is then used to guide the Transformer in denoiser. Notably, since the distribution of the latent space with the size of $\mathbb{R}^{N\times C}$ (e.g, $16\times 256$) is much simpler than that of images with size $\mathbb{R}^{H\times W\times N_{\lambda}}$, the prior feature can be generated with a small number of iterations $T$, corresponding to paper [53].
@@
$\nabla_{\bm{\theta}}\left\|\bm{\epsilon}-\bm{\epsilon}_{\theta}\left(\sqrt{\bar{\alpha}_{t}}\bm{z}_{GT}+\sqrt{1-\bar{\alpha}_{t}}\bm{\epsilon},{\bm{c}},t\right)\right\|_{2}^{2},$ (11)
where $\bm{z}_{GT}$ and ${\bm{c}}$ are ground-truth prior features and the latent condition representations defined above; $t\in[1,T]$ is the randomly sampled time step; $\bm{\epsilon}\sim\mathcal{N}(0,\mathbf{I})$ denotes the sampled Gaussian noise. We employ DDPM [22] for diffusion as the original LDM [53]. Considering small prior size and time efficiency, we adopt a simple denoising network consisting of several MLP layers for fast diffusion denoising. All the parameters are jointly updated in the network with the objective loss function of the second phase, including:
the DUN, the feature encoder $\text{LE}^{\prime}$, and the diffusion denoising network $\bm{\epsilon}_{\theta}$. The objective loss function of the second phase can be stated as:
@@
Here, we do not use Eq. (11) as $\mathcal{L}_{\text{diff}}$ because it only trains diffusion at ‘t’-th step while we execute the all-time-step together and let the LDM directly predict $\hat{\bm{z}}$. The entire two-phase training procedure is summarized in Algorithm 1.
@@ 4.3 Aggregate Features by Trident Transformer
Figure 4: (a) The Trident Transformer in Fig. 3(d). (b)-(d) are the detailed sub-modules. ${{\bf U}}_{i}$ is the input feature. The prior feature ${\bf Z}_{i}$ is sent into the prior flow.
@@
Previous HSI reconstruction methods usually only exploit the relation between spatial and spectral, both externally and internally. However, the spatial-spectral relations are challenging to explore only with compressed measurements. Therefore, we design a Transformer, named Trident Transformer (TT), to effectively aggregate high-quality degradation-free prior knowledge for compensation.
@@
Firstly, inspired by the multi-scale operations in previous papers [12, 68] with hierarchical structures, we downsample the prior to obtain the multi-scale prior representations along with the U-shape levels in Fig. 3(d). Specifically, three downsampling layers are employed, and the outputs contain prior features of three scales, stated as:
$\bm{z}^{i}=\left\{\begin{array}[]{ccl}\bm{z}_{GT}\text{~{}or~{}}\hat{\bm{z}},&{if}&i=1,\\ \text{downsample}(\bm{z}^{i-1}),&{if}&i>1\end{array}\right.~{},$ (13)
@@
where $\bm{z}^{i}\in\mathbb{R}^{\frac{N}{2^{i}-1}\times 2^{i-1}C},~{}i=1,2,3$. For Phase I, $\bm{z}^{1}=\bm{z}_{GT}$, which is computed in the first phase training; For Phase II, $\bm{z}^{1}=\hat{\bm{z}}$, which is utilized for training and inference in the second phase.
@@
As shown in Fig. 4, our Trident Transformer includes three branches: spatial flow, cross-spectral flow, and cross-prior flow. Each branch shares the information flow with others and is then fused by the aggregation layer and a feed-forward network (FFN). Before the embedding layer, the input feature at $i$-th scale ${{\bf U}}_{i}\in\mathbb{R}^{H_{i}\times W_{i}\times C_{i}}$ is split into ${{\bf U}}^{C}_{i}\in\mathbb{R}^{H_{i}\times W_{i}\times\frac{C_{i}}{2}}$ and ${{\bf U}}^{S}_{i}\in\mathbb{R}^{H_{i}\times W_{i}\times\frac{C_{i}}{2}}$ along the channel dimension, denoting cross flow input and spatial flow input respectively. The spatial flow consists of a series of MBlocks without batch norms.
@@
Cross Spectral Flow In the cross spectral flow (CSF) module, as shown in Fig. 4 (c), we design asymmetric cross-scale multi-head self-attention (ACS-MHSA). Pansharpening (PAN) is a technique of using a high-resolution (HR) panchromatic and a lower-resolution (LR) HSI to generate an HR-HSI. Compared with directly capturing HR-HSI, it requires less amount of data. Inspired by the PAN, this flow primarily focuses on the spectral dimension and aims to save computational burden according to the spatial size. Specifically, we compress the spatial resolution of the query embedding (${{\bf Q}}$) and key embedding (${{\bf K}}$) to $\frac{1}{4}$ and expand its channel twice. Considering that CASSI measurements are shifted along one axis, there are more spatial correlations along this axis. Thus, after establishing the spectral correlation, we use an asymmetric dilation convolution (DConv) with kernel size $3\times 5$ on the value embedding (${{\bf V}}$) to obtain larger perceptual field information along the shifted axis with expanded channel dimension and unchanged spatial dimension. Embedding ${{\bf Q}}$, ${{\bf K}}$, ${{\bf V}}$, and spatial compensation ${{\bf P}}_{i}^{S}$ from spatial flow are formulated as:
${{\bf Q}}_{i}^{CS}$ $={{\bf W}}^{QCS}{{\bf U}}_{i},$ ${{\bf K}}_{i}^{CS}$ $={{\bf W}}^{KCS}{{\bf U}}_{i},$ ${{\bf V}}_{i}^{CS}$ $={{\bf W}}^{VCS}{{\bf U}}_{i},$ (14)
${{\bf P}}_{i}^{SQK}$ $=\downarrow({{\bf W}}^{PSQ}{{\bf Q}}_{i}^{S}),$ ${{\bf P}}_{i}^{SV}$ $={{\bf W}}^{PSV}{{\bf Q}}_{i}^{S},$ ${{\bf M}}_{i}$ $={{\bf Q}}_{i}^{CS}({{\bf K}}_{i}^{CS})^{\top},$ (15)
where ${{\bf W}}^{*}$ represents the weights of bias-free convolution, and $\downarrow$ is downsampling. The procedure of asymmetric cross-scale self-attention can be formulated as:
[…]
@@
Cross Prior Flow Cross prior flow (CPF) in Fig. 4 (d) is a variable shared multi-head cross-attention. The query in this flow is borrowed from the value of CSF which is extracted from a large perceptive field with more spatial information. In this way, the prior could facilitate to compensate for spatial deficiency. Compared to the spectral recovery, the spatial recovery is typically more challenging. Our manipulation can be formulated as:
${MHSA_{i}^{CP}}({{\bf U}}_{i})=W_{c1}{{\bf V}}\odot\text{Softmax}({{\bf K}}\odot{{\bf Q}}/\alpha),$ (17)
@@
Flow Interaction and Aggregation In order to compensate for the deficiency of spatial information in CSF and CPF, and the spectral information in the spatial flow, we fuse the compensation information together to reconstruct hyperspectra images. As shown in Fig. 4, the colorful arrows represent the information interactions between each flow. Specifically, information of each module modulates with other flows, where the $1\times 1$ convolutions serve as compensation bridges. The aggregation part consists of concatenation, convolution layers, and an activation function (details can be seen in SM) for a weighted combination of each flow output. In our Trident Transformer, the prior knowledge learned from the clean images will provide compensation for reconstruction in both spatial and spectral details, avoiding the influence of degraded measurements.
Algorithm 1 Two-phase Training Strategy
1:Dataset $\mathcal{D}=\{(\bm{x}^{(n)},\bm{y}^{(n)}\}^{N}_{n=1}$; Sensing matrix $\bm{A}$; Random initialized parameter of DUN $\phi_{DUN}$, latent encoder LE network $\phi_{LE}$, conditional encoder LE’ network $\phi_{CLE}$, diffusion denoising network $\phi_{\epsilon}$;
2:while Not Converge do $\triangleright$ Phase I training, $\leftarrow$’ denotes update
3:     $\bm{z}_{GT}\leftarrow\text{LE}({\bm{I}}_{\text{E}}|\phi_{LE})$; $\hat{x}\leftarrow\text{DUN}(\bm{y},\bm{A},\bm{z}_{GT}|\phi_{DUN})$
4:     Jointly update $\phi_{DUN}\text{ and }\phi_{LE}$ by $\mathcal{L}_{rec}$;
5:Freeze $\phi_{LE}$;
6:while Not Converge do $\triangleright$ Phase II training
7:     ${\bm{c}}\leftarrow\text{LE}^{\prime}(\bm{y}_{norm}|\phi_{CLE})$; $\hat{\bm{z}}\leftarrow\text{Diff}({\bm{c}}|\phi_{\epsilon})$;
8:     $\bm{z}_{GT}\leftarrow\text{LE}({\bm{I}}_{\text{E}}|\phi_{LE})$; $\hat{x}\leftarrow\text{DUN}(\bm{y},\bm{A},\hat{\bm{z}})|\phi_{DUN})$;
9:     Jointly update $\phi_{DUN},\phi_{CLE},\text{ and }\phi_{\epsilon}$ by $\mathcal{L}_{all}$ in Eq. (12);
Table 1: The results of PSNR in dB (top entry in each cell), SSIM (bottom entry in each cell) on the 10 synthetic spectral scenes.‘-3stg’ denotes the network with 3 unfolding stages. ‘Avg’ represents the average of 10 scenes. Bold: Best.
@@
38.08 41.84 43.77 47.99 37.97 38.30 38.82 36.15 42.53 35.48 40.09
Ours-9stg 0.969 0.982 0.983 0.993 0.980 0.980 0.973 0.979 0.984 0.970 0.979
38.08 41.85 43.83 48.04 38.00 38.32 38.94 36.20 42.81 35.54 40.16
Ours-10stg 0.970 0.984 0.984 0.993 0.982 0.982 0.974 0.979 0.984 0.970 0.980
Table 2: Performance and computational efficiency comparisons with recent methods.
Method Venue PSNR (dB) Params (M) FLOPs (G) Infer. Time (ms) Training Time (h)
PADUT-12stg ICCV’23 38.89 5.38 90.46 749.94 123.3
RDLUF-9stg CVPR’23 39.57 1.89 231.09 913.34 155.5
Ours-9stg - 40.09 2.78 88.68 1096.58 143.6
Table 3: Ablation study of different modules in our 3-stage unfolding network. ‘PSNR’ is the average of 10 synthetic scenes. ‘FLOPs (G)’ denotes the FLOPs in testing.
Method Projection SF CSF Atten. CPF LE Input Diffusion PSNR FLOPs (G)
w/o GC-GAP Basic GAP ACS ${\bm{I}}_{\text{E}}$ 38.01 29.84
w/o spatial flow GC-GAP ACS ${\bm{I}}_{\text{E}}$ 37.66 30.13
Basic MHSA GC-GAP Basic ${\bm{I}}_{\text{E}}$ 38.31 36.80
w/o prior flow GC-GAP ACS ${\bm{I}}_{\text{E}}$ 37.89 30.37
Inaccurate guidance GC-GAP ACS $\bm{y}$ 37.63 33.80
w/o diffusion GC-GAP ACS ${\bm{I}}_{\text{E}}$ 37.61 33.70
Our full model GC-GAP ACS ${\bm{I}}_{\text{E}}$ 38.40 33.80
Figure 6: The real data visual comparisons. 2 out of 28 wavelengths are selected.